OpenStudy (anonymous):

divide / simplify z^2-25/z DIVIDE BY z-5/z+6

OpenStudy (saifoo.khan):

\[\frac{z^2-25}{z} \times \frac{z+6}{z-5}\]

OpenStudy (saifoo.khan):

\[\frac{(z-5)(z+5)}{z} \times \frac{z+6}{z-5}\]

OpenStudy (saifoo.khan):

\[\frac{\cancel{(z-5)}(z+5)}{z} \times \frac{z+6}{\cancel{z-5}}\]

OpenStudy (anonymous):

z^2 - 25 z + 6 (z - 5) (z + 5) (z + 6) z^2 + 11z + 30 -------- * ------- = ------------------- = --------------- z z - 5 (z - 5) * z z

OpenStudy (anonymous):

so the then i cross multiply so it would the be (z+5)(z-6) right?

OpenStudy (saifoo.khan):

over "z" as well.

OpenStudy (anonymous):

I hv solved it fully......